OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today? 330 510 520 660

OpenStudy (australopithecus):

x + y =840 3(x) - 2(y) = 870

OpenStudy (australopithecus):

x is the one step equations y is the two step equations

OpenStudy (australopithecus):

do you follow how I got those equations? The trick is just to write down the equation as you read

OpenStudy (australopithecus):

I assume you can solve from here

OpenStudy (anonymous):

i still need help.. i am really not good at word problems.

OpenStudy (australopithecus):

well do you understand how I got those equations?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

do you know how to use system of equations?

OpenStudy (anonymous):

yes i believe so

OpenStudy (australopithecus):

then you can solve this problem

OpenStudy (anonymous):

i know, but i dont know exactly i what i could do to get the correct answer

OpenStudy (australopithecus):

how many one-step equations auditioned today? x is the number of one step equations so solve for y x + y =840 y = 840 - x sub into the second equation 3(x) - 2(y) = 870 3(x) - 2(840 - x) = 870 solve for x

OpenStudy (australopithecus):

follow?

OpenStudy (australopithecus):

when using the system of equations you are simply just finding where two lines intersect.

OpenStudy (anonymous):

would it be 3x - 1680 + 2x?

OpenStudy (anonymous):

when getting rid of the parenthesis?

OpenStudy (australopithecus):

looks right

OpenStudy (australopithecus):

I HAVE THE POWER!

OpenStudy (anonymous):

would i subtract 2 from 3 or 3 from 2?

OpenStudy (australopithecus):

3(x) - 2(840 - x) = 870 3x + (-1)(2)(840 - x) = 870 3x + (-2)(840-x) = 870 3x + (-2)(840) + (-x)(-2) = 870 3x + (-2)(840) + 2x = 870 5x + (-2)(840) = 870 5x +(-2)(840) - (-2)(840) = 870 - (-2)(840) 5x = 870 - (-2)(840) \[\frac{5x}{5} = \frac{870-(-2)(840)}{5}\] \[x = \frac{870-(-2)(840)}{5}\]

OpenStudy (anonymous):

how would i divide this. sorry i am not good enough for algebra 1

OpenStudy (australopithecus):

870 - (-2)(840) is just a number I was too lazy to multiply it out 870 + (-1)(-2)(840) 870 + (2)(840) 870 + 1680 2550

OpenStudy (australopithecus):

follow?

OpenStudy (australopithecus):

2550/5 = x

OpenStudy (australopithecus):

use a calculator if you are allowed if not then just

OpenStudy (australopithecus):

do it the long way

OpenStudy (anonymous):

510?

OpenStudy (australopithecus):

yup

OpenStudy (anonymous):

thank you

OpenStudy (australopithecus):

yup